OpenStudy (anonymous):

PLEASE HELP How do you factor 10Sin^2+3Sin-4=0

OpenStudy (anonymous):

What comes after sin² ?

OpenStudy (anonymous):

+3sin-4

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

10Sin(theta)^2+3sin(theta)-4=0

OpenStudy (anonymous):

So i think we must factor, but i can't find one that works

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

Try this: Let \(\sin \theta\) be \(x\) then factor.

OpenStudy (anonymous):

yeah, thats what my teacher taught me. But i can't find how to factor this normally

OpenStudy (anonymous):

You can factor this normally. HINT: (2x _ ___)(5x _ ___)

OpenStudy (anonymous):

Oh ok, i'll try that. One sec. at first i tried 10x, x

OpenStudy (anonymous):

We can use -1, and +4

OpenStudy (anonymous):

so then theta would be +1/2 and -4/5

OpenStudy (anonymous):

i just wrote down the question i remember. Would it ask what the exact values of <theta>360?

OpenStudy (anonymous):

Would i have 4 answers or something?